Diffence between two loans is $10000, both loans have 5% annual interest rate. If the total interest is $2000 after one year how much was each loan?
Answer by Fermat(127)

Call the two loans L1 and L2. with L1 > L2
Then, L1 - L2 = 10,000
commom interest rate = 5%
I1 = interest on L1
I2 = interest on L2
I1 = %5*L1 = 0.05L1
I2 = %5*L2 = 0.05L2
I1 + I2 = 2000
.: 0.05L1 + 0.05L2 = 2000
0.05(L1 + L2) = 2000
dividing both sides by 0.05
L1 + L2 = 40,000 ----------------(1)
but
L1 - L2 = 10,000 ----------------(2)
adding (1) and (2),
2L1 = 50,000
L1 = 25,000
===========
L2 = 15,000
===========